OpenStudy (anonymous):

Guys can you pls help me answering this prob. Evaluate lim x-2 over x^2-1 as x approaches 2?

OpenStudy (anonymous):

replace x by 2, get zero

OpenStudy (anonymous):

\[\lim_{x\rightarrow 2}\frac{x-2}{x^2-1}=\frac{2-2}{2^2-1}=\frac{0}{3}=0\]

OpenStudy (anonymous):

@ Sattelite thanks for me helping me. :)

OpenStudy (anonymous):

yw most every function given by and equation is continuous on its domain, meaning to find a limit you can simply evaluate the funtion. you will only have more work to do the funciton is undefined at the limit point, specifically if you get \[\frac{0}{0}\] when you evaluate the function

OpenStudy (anonymous):

could you help me again answering my take home quiz in calculus? I have 19 questions left unanswered. :(

OpenStudy (anonymous):

post i bet you will get lots of answers better to post them one at a time

OpenStudy (anonymous):

yes i will thank you so much :)

OpenStudy (anonymous):

2nd question: Find lim (sin2xcot4x) X->0.